# Feature Importance-Aware Deep Joint Source-Channel Coding for Computationally Efficient and Adjustable Image Transmission

### v1

@@ V Conclusion
In this paper, we introduced FAJSCC, a novel approach for efficient image transmission with adjustable computational resource allocation. Unlike existing deepJSCC research that uniformly allocates equal processing power across all features, FAJSCC dynamically allocates resources based on feature importance. Specifically, self-attention is applied to important features, while spatial attention is used for less important features, optimizing computational efficiency. Due to this importance-wise feature computation, FAJSCC achieves superior performance compared to baseline models while consuming fewer computational and memory resources than SwinJSCC. Furthermore, FAJSCC introduces flexible computational resource control, allowing independent adjustments to the encoder and decoder’s computational resources without requiring knowledge of each other’s computational budget. This adjustable property ensures high performance across different computational constraints, outperforming SwinJSCC under varying computational resource availability. By effectively balancing feature importance, computational efficiency, and reconstruction performance, FAJSCC becomes the first deepJSCC model capable of importance-aware feature processing and computational resource control with high performance across various computational resources and communication channel conditions. Moreover, our deepJSCC performance analysis implies that the decoder’s error correction function consumes the largest fraction of computational complexity. A promising future direction is to develop an improved FAJSCC that dynamically allocates computational resources to the decoder’s error correction module based on channel conditions while minimizing unnecessary computation on the encoder side. This would maximize deepJSCC’s efficiency and adaptability in real-world communication systems.

### v2

@@ V Conclusion
In this paper, we introduced FAJSCC, a novel framework for efficient image transmission with adjustable computational complexity based on the attention family tree. Leveraging axis-dimension specialized computation, FAJSCC reduces computational cost while preserving feature expressiveness. Building on this foundation, our selective deformable self-attention mechanism further enhances transmission performance by adaptively refining the most important features. In addition, our attention family tree efficiently provides attention features for axis-specialized computation and selective deformable self-attention by eliminating overlapping computations. The flexibility of selective deformable self-attention ensures strong performance under varying computational constraints, consistently outperforming SwinJSCC. Moreover, FAJSCC is the first deepJSCC model capable of independently adjusting the computational resources of the encoder and decoder. Our performance analysis reveals that the decoder’s noisy received signal perception function consumes the largest fraction of computational complexity. A promising future direction is to develop an improved FAJSCC that dynamically allocates computational resources to the decoder’s signal perception module based on channel conditions while minimizing unnecessary computation on the encoder side. This would maximize deepJSCC’s efficiency and adaptability in real-world communication systems.